OpenStudy (anonymous):

Need help with slope. will give medal and fan Select the proper equation in point-slope form for the given point and slope. Point: (–3, 7); Slope: 4

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Luigi0210

OpenStudy (danjs):

y - y1 = m(x - x1) point (x1,y1) slope = m

OpenStudy (anonymous):

y - 7 = 4(x - 3) y - 7 = 4(x + 3) y + 7 = 4(x + 3) y + 7 = 4x - 12

OpenStudy (anonymous):

those are my options

OpenStudy (danjs):

So using the last post i put up, just put the numbers in, which one matches?

OpenStudy (anonymous):

A?

OpenStudy (danjs):

close, y - 7 + 4(x - (-3))

OpenStudy (danjs):

minus a negative will end up in x+3

OpenStudy (anonymous):

so B.

OpenStudy (danjs):

yep

OpenStudy (anonymous):

can you help with a few more?

OpenStudy (danjs):

ok

OpenStudy (anonymous):

Select the proper equation in point-slope form for the given point and slope. Point: (1, -7); Slope: (-2/3)

OpenStudy (danjs):

It is the exact same thing, what do you think first?

OpenStudy (danjs):

y - y1 = m(x - x1)

OpenStudy (anonymous):

ok ill check

OpenStudy (danjs):

m = -2/3 x1 = 1 y1 = -7 remember minus a negative leaves a Plus

OpenStudy (anonymous):

so its y - 7 = (-2/3) (x+1)

OpenStudy (anonymous):

is it right?

OpenStudy (danjs):

You got a little bit confused with the signs. Put the numbers in first then deal with the signs. (y -(-7)) = -2/3 ( x - 1)

OpenStudy (anonymous):

oh ok but other than that its right?

OpenStudy (anonymous):

Identify the point (x1, y1) from the equation: y - 9 = 3(x + 2) (-2, 9) (9, -2) (2, -9) (-8, 2)

OpenStudy (danjs):

(y+7) = -2/3 (x - 1)

OpenStudy (danjs):

The next question is again, the exact same question, the equation is in the form y - y1 = m(x - x1) ---that is all you need to remember so what is x1 and y1? (x1 , y1)

OpenStudy (danjs):

remember a plus sign is the same as minus a negative for example (x + 3) = (x -(-3))

OpenStudy (anonymous):

its C

OpenStudy (danjs):

close again, the general form is y - y1 = m(x - x1) y - 9 = 3(x-(-2))

OpenStudy (danjs):

x1 = -2 y1 = +9

OpenStudy (anonymous):

so its A. alright thank you for your help.

OpenStudy (danjs):

no prob